OpenStudy (anonymous):

If the measure of each exterior angle of a regular polygon is 18°, determine how many sides the polygon has. A. 20 sides B. 22 sides C. 18 sides D. 16 sides

OpenStudy (anonymous):

Try solving for n in the formula E=360/n

OpenStudy (anonymous):

Measure of Interior Angle = (n-2) × 180° / n Interior Angle = 180 - Exterior Angle Thus, 180 - Exterior Angle = (n-2) x 180 / n 180 - 18 = (n-2)x180 / n and solve for n.

OpenStudy (anonymous):

20 degrees ?

OpenStudy (anonymous):

yes because it becomes n=360/E and E=18 so n=20.

OpenStudy (anonymous):

thanks